OpenStudy (anonymous):

I am not doing this trig right given tan(A+B) =4; tanB=3/5 using formula (tanA + tanB)/(1-tanA*tanB) so replacing tanB with 3/5 (tanA +3/5)/(1-tanA*3/5)=4 I am at a total block can you help please?

OpenStudy (amistre64):

the rest is basic algebra ... multiply, distribute, collect like terms ... factor, etc

OpenStudy (amistre64):

(x+3)/(1-3x) = 4 how do you solve for x?

OpenStudy (anonymous):

sorry my computer locked up

OpenStudy (anonymous):

I got tanA+3/5=4(1-tanA*(3/5)) tanA+3/5=4-4*tanA*(3/5) tanA+4tanA*3/5=4-(3/5) tanA+(12tanA/5)=17/5 (5tanA/5)+(12tanA/5)=17/5 17tanA/5=17/5 tanA=1 Thanks for the push!